OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) can someone please help me solve this inequaltiy???

OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) distribute the -2 to the (x-3), then the - sign to the (x+3) -2x+6 \ge \] 5-x-3 combine like terms \[-2x+6 \ge -x+2\] get x alone, and solve from there.

OpenStudy (anonymous):

\[-2x+6 \ge -5x-15\]